At the end of 2017, Alphabet, Inc., reported stockholders’ equity of $152,502 million and total assets of $197,295 million. Its balance in stockholders’ equity at the end of 2016 was $139,036 million. Net income in 2017 was $12,662 million.

a. Calculate Alphabet, Inc., return on equity ratio for 2017. (Round your answer to one decimal place.)

b. Calculate its debt-to-equity ratio as of December 31, 2017. (Hint: Apply the accounting equation to determine total liabilities.) Round your answer to one decimal place.
Answer #1

Return on equity = Net income / Equity

Net income = $12,662

Equity = $152,502

= (12,662/152,502)*100

= 8.3%

B)

Debt equity ratio = Debt / Equity

Debt = Total assets - Equity

= 197,295 - 152,502

= $44,793

Ratio

= 44,793/152,502

= 0.294

=29.4%